OpenStudy (anonymous):

Two tiny, spherical water drops, with identical charges of -2.80 10-17 C, have a center-to-center separation of 2.0 cm. (a) What is the magnitude of the electrostatic force acting between them? N (b) How many excess electrons are on each drop, giving it its charge imbalance? electrons

OpenStudy (johnweldon1993):

The force acting between them would be \[\large F = k\frac{q_1\times q_2}{r^2}\] However, here \(\large q_1 = q_2\) so lets just write \[\large F = k\frac{q^2}{r^2}\] So for part A we would have \[\large F = 9\times 10^9 \frac{Nm^2}{C^2}\times \frac{(-2.80\times 10^{-17}C)^2}{(0.02m)^2} = ?\] Part B, We can use the expression \(\large ne = q\) where e = the charge of an electron and q is the net charge So we would solve that for 'n' using the fact that we are given the net charge in the question \[\large n = \frac{q}{e}\] \[\large n = \frac{-2.80\times 10^{-17}C}{-1.6\times 10^{-19}C}=?\]